# QUANTUM MECHANICS OF THE NONRELATIVISTIC YANG MODEL

### v1

@@ Abstract
To give an interpretation to the physics associated to the Yang algebra, it is natural to identify $\alpha$ and $\beta$ with the cosmological constant and the inverse of the Planck mass squared, respectively. The parameters are therefore exceedingly small for experimental scales related to particle physics and gravity. Although from a mathematical point of view all values of positions and momenta are allowed (except that for positve values of the coupling constants $x^{2}<1/\alpha$ and $p^{2}<1/\beta$), the theory is supposed to be reliable only for values of $x^{2}\ll 1/\alpha$ and $p^{2}\ll 1/\beta$.
Although the Yang model has attracted growing interest recently, only its formal properties have been investigated [6], with the exception of ref. [7], where some implications on statistical mechanics and on the uncertainty principle were considered.
@@
In this framework, we can discuss the modification of the uncertainty relation induced by the deformed symplectic structure. Deformed uncertainty relation have earned much interest in the last years, since they may emerge in a variety of situations, including string theory, quantum gravity, noncommutative geometry or gedanken experiments [8]. However, also the curvature of spacetime leads to a deformation of the uncertainty relation [9]. Both deformation can be unified in the so called Extended Generalized Uncertainty Principle (EGUP) [10]. Of course, as we shall show, it is this last case that occurs in the Yang model.
In the context of deformed uncertainty relations, also the investigation of simple quantum mechanical models, like the free particle or the harmonic oscillator, is interesting and has been performed by several authors [11]. In particular, it has been shown that in some cases the free particle has discrete spectrum, while the harmonic oscillator always displays corrections with respect to the standard case.
@@
In the present case the investigation of the deformations of the uncertainty relation and of the dynamics can be done only if one considers a specific realization of the model on phase space, since, as noted above, it is otherwise difficult to give a physical interpretation to the scalar generator of the algebra. Moreover, to avoid computational complications, we shall limit our study to the one-dimensional case and to the leading order in $\hbar$. We shall consider both positive and negative values of the coupling parameters, since they lead to very different physical predictions.
@@
${\hat{x}}_{\mu}=x_{\mu}\sqrt{\sqrt{1-\alpha\beta(x\!\cdot\!p)^{2}}-\beta p^{2}}\,\qquad{\hat{p}}_{\mu}=p_{\mu}\sqrt{\sqrt{1-\alpha\beta(x\!\cdot\!p)^{2}}-\alpha x^{2}}\ ,$ (3)3( 3 )
@@
$K=\sqrt{1-\alpha{\hat{x}}_{\rho}^{2}-\beta{\hat{p}}_{\rho}^{2}-{\alpha\beta\over 2}M_{\rho\sigma}^{2}}=\sqrt{1-\alpha{\hat{x}}_{\rho}^{2}-\beta{\hat{p}}_{\rho}^{2}-\alpha\beta\,{{\hat{x}}_{\rho}^{2}{\hat{p}}_{\sigma}^{2}-({\hat{x}}_{\rho}{\hat{p}}_{\rho})^{2}\over K^{2}}}.$ (5)5( 5 )
From (3) it follows that if $\alpha\beta>0$, the scalar product $x\!\cdot\!p$ must have an upper bound $1/\sqrt{\alpha\beta}$, while this is not true if $\alpha\beta<0$. Moreover, if $\alpha$ or $\beta$ are positive, one must require $x^{2}\leq 1/\alpha$ or $p^{2}\leq 1/\beta$, respectively.
@@
In the nonrelativistic model, the commutation relations are the same as (1), but now the indices run from 1 to 3, $\eta_{\mu\nu}\to\delta_{\mu\nu}$ and the algebra reduces to $o(5)$.
3. Generalized uncertainty principle A discussion of the properties of the Yang model in the realization of the previous section is however extremely difficult, even in the 3D Euclidean version. The problem is however highly simplified in the one-dimensional limit. Although in one dimension one cannot speak of noncommutativity and curvature of spacetime, this limit has nevertheless proven to maintain the most interesting features of its higher-dimensional counterparts in similar models [5,13], since its behaviour essentially coincides with that of the higher-dimensional radial equation, corresponding to the zero angular momentum sector. In the Snyder case this limit has also been widely investigated in the context of minimal length models [11]. Recently, the appearance of deformed uncertainty relation in the 1D Yang model has independently been discussed in [7].
In one dimension, the commutation relation for ${\hat{x}}$ and ${\hat{p}}$ in our approximation with $K$ given by (5), simplifies to
$[{\hat{x}},{\hat{p}}]=i\hbar\sqrt{1-\alpha{\hat{x}}^{2}-\beta{\hat{p}}^{2}}$ (7)7( 7 )
It follows that for $\alpha,\beta>0$, the acceptable values of ${\hat{x}}^{2}$ are bounded by $1/\alpha$, and those of ${\hat{p}}^{2}$ by $1/\beta$, and in general $\alpha{\hat{x}}^{2}+\beta{\hat{p}}^{2}\leq 1$. No such bound is instead present for negative values of the coupling constants.
Assuming for simplicity $\langle{\hat{x}}\rangle=\langle{\hat{p}}\rangle=0$ the commutation relations (7) imply the generalized uncertainty relation, analogous to EGUP,
$\Delta{\hat{x}}\Delta{\hat{p}}\geq{\hbar\over 2}\left|\langle[{\hat{x}},{\hat{p}}]\rangle\right|={\hbar\over 2}\sqrt{1-\alpha(\Delta{\hat{x}})^{2}-\beta(\Delta{\hat{p}}^{2})}$ (8)8( 8 )
@@
$(\Delta{\hat{x}})\geq\hbar\sqrt{{1-\beta(\Delta{\hat{p}})^{2}\over\hbar^{2}\alpha+4(\Delta{\hat{p}})^{2}}},\qquad(\Delta{\hat{p}})\geq\hbar\sqrt{{1-\alpha(\Delta{\hat{x}})^{2}\over\hbar^{2}\beta+4(\Delta{\hat{x}})^{2}}}$ (9)9( 9 )
The behaviour of $\Delta x$ in function of $\Delta p$ in the various cases is reported in fig. 1:
@@
For $\alpha,\beta>0$, the uncertainties satisfy the bounds $0\leq\Delta p\leq 1/\sqrt{\beta}$, and $0\leq\Delta x\leq 1/\sqrt{\alpha}$, as is obvious because of the bounds imposed on $x$ and $p$.
For $\alpha,\beta<0$, $\Delta p\geq\hbar\sqrt{|\alpha|}/2$, and $\Delta x\geq\hbar\sqrt{|\beta|}/2$, hence both position and momentum have a minimal uncertainty independent of that of the conjugated variable.
A more complicated behaviour occurs if $\alpha$ and $\beta$ have different signs. If $\alpha<0$ and $\beta>0$, $\Delta p$ must satisfy the bounds $\hbar\sqrt{|\alpha|}/2\geq\Delta p\geq 1/\sqrt{\beta}$, while $\Delta x$ can take any positive value.
Finally, If $\alpha>0$ and $\beta<0$, $\Delta p$ can take any positive value, while $\hbar\sqrt{|\beta|}/2\geq\Delta x\geq 1/\sqrt{\alpha}$.
It is also clear that if either $\alpha$ or $\beta$ is positive, it is possible to have states with vanishing $\Delta x\Delta p$, while this cannot occur if both $\alpha$ and $\beta$ are negative. It seems therefore that only the case of positive $\alpha$ and $\beta$ can satisfy the standard uncertainty bound. However, also for negative values of the coupling constant states with vanishing uncertainty are achieved only for values of $\Delta x$ and $\Delta p$ close to their limit values, which are assumed to be out of the range of applicability of the theory.
@@
$\Delta x\Delta p\geq{\hbar\over 2}\,{1+\alpha\Delta x^{2}+\beta\Delta p^{2}\over 1+\hbar\sqrt{\alpha\beta}}$ (10)10( 10 )
@@
$\Delta p_{min}={\hbar\sqrt{\alpha}\over\sqrt{1+2\hbar\sqrt{\alpha\beta}}},\qquad\Delta x_{min}={\hbar\sqrt{\beta}\over\sqrt{1+2\hbar\sqrt{\alpha\beta}}}$ (11)11( 11 )
@@ 4. One-dimensional dynamics
Clearly, the quantum dynamics of the Yang model differs from the standard one, because of the deformation of the commutation relations. In one dimension, the dynamics of the free particle and of the harmonic oscillator has been studied for the Snyder model (also in the context of the generalized uncertainty principle) [11,5] and TSR [13], displaying a discrete spectrum for the free particle with positive coupling constants, and a deformation of the energy spectrum of the harmonic oscillator.
Let us first consider the solution of the one-dimensional Schrödinger equation for a free particle in the Yang model. The equation is not exactly solvable, so we make an expansion in the parameters $\alpha$ and $\beta$. In one dimension, eqs. (3) become at leading order
${\hat{x}}_{\mu}=x_{\mu}-{\beta\over 4}(x_{\mu}p^{2}+p^{2}x_{\mu}),\qquad{\hat{p}}_{\mu}=p_{\mu}-{\alpha\over 4}(p_{\mu}x^{2}+x^{2}p_{\mu})$ (12)12( 12 )
where we have chosen a Hermitian realization. For simplicity we consider the case $\alpha,\beta$ positive. One can now work either in position or momentum representation. In the first case,
$x_{\mu}\to x_{\mu},\quad p_{\mu}\to-i{\partial\over\partial x_{\mu}}$ (13)13( 13 )
and the Schrödinger equation for a free particle becomes
$({\hat{p}}^{2}-2mE)\,\psi=-(1-\alpha x^{2})\,{d^{2}\psi\over dx^{2}}+2\alpha x{d\psi\over dx}+\left({\alpha\over 2}-2mE\right)\psi=0$ (14)14( 14 )
Imposing the boundary conditions $\psi(-1/\sqrt{\alpha})=\psi(1/\sqrt{\alpha})=0$, the solution is
$P_{n}(\sqrt{\alpha}\,x),$ (15)15( 15 )
@@
$E={\alpha\over 2m}\left(n^{2}+n+{1\over 2}\right)$ (16)16( 16 )
@@
$p_{\mu}\to p_{\mu},\quad x_{\mu}\to i{\partial\over\partial p_{\mu}}$ (17)17( 17 )
@@
$({\hat{p}}^{2}-2mE)\,\psi=\alpha p^{2}{d^{2}\psi\over dp^{2}}+2\alpha p\,{d\psi\over dp}+\left(p^{2}-2mE+{\alpha\over 2}\right)\psi=0$ (18)18( 18 )
The solution is given by
$\psi=j_{n}\left({p\over\sqrt{\alpha}}\right)$ (19)19( 19 )
with $j_{n}$ a spherical Bessel function with index such that $n(n+1)={2mE\over\alpha}-{1\over 2}$, recovering the previous result.
@@
$H={{\hat{p}}^{2}\over 2m}+{m\omega^{2}{\hat{x}}^{2}\over 2}$ (20)20( 20 )
@@
$\Big{(}1-A^{2}x^{2}\Big{)}\,{d^{2}\psi\over dx^{2}}-2A^{2}x{d\psi\over dx}+\left(2mE-{A^{2}\over 2}-m^{2}\omega^{2}x^{2}\right)\psi=0$ (21)21( 21 )
@@
$A^{2}=\alpha+\beta m^{2}\omega^{2}$ (22)22( 22 )
@@
$(1-z^{2})\,{d^{2}\psi\over dz^{2}}-2z^{2}{d\psi\over dx}+\left({2mE\over A^{2}}-{1\over 2}-{m^{2}\omega^{2}\over A^{4}}z^{2}\right)\psi=0$ (23)23( 23 )
@@
$\psi_{n}=S_{0n}(Ax)$ (24)24( 24 )
Notice that eq. (23) has not a smooth limit for $A\to 0$. From (23) it follows that
$E_{n}={A^{2}\over 2m}\left(\lambda_{n}-{1\over 2}\right)$ (25)25( 25 )
@@
$\lambda_{n}=\sum_{k=0}^{\infty}l_{2k}\gamma^{2k},\qquad{\rm with}\quad\gamma^{2}={m^{2}\omega^{2}\over A^{4}}$ (26)26( 26 )
@@
$l_{0}=n(n+1),\qquad l_{2}={1\over 2}\left(1-{1\over(2n-1)(2n+3)}\right),\qquad{\rm etc.}$ (27)27( 27 )
@@
$\lambda_{n}=2\left(n+{1\over 2}\right)\gamma-{1\over 2}\left(n^{2}+n+{3\over 2}\right)+o\left({1\over\gamma}\right)$ (28)28( 28 )
@@
$E_{n}=\left(n+{1\over 2}\right)\omega-{A^{2}\over 4m}\left(n^{2}+n+{3\over 2}\right)+o\left({A^{4}\over m^{2}\omega}\right)$ (29)29( 29 )
Besides the standard term, one has corrections proportional to powers of $A^{2}$. Also this result is analogous to TSR [13], where $E_{n}=\left(n+{1\over 2}\right)\omega+{A^{2}\over 2m}\left(n^{2}+n+{1\over 2}\right)+\dots$, but the coefficients of the correction terms are different, and in particular have opposite sign.
@@
5. Final remarks We have considered the quantum mechanics of the one-dimensional Yang model in its simplest realization on phase space. The investigation of this simple case is relevant because, as shown in similar situations [11,5], the results are qualitatively similar to those valid in 3 dimensions. The results can be compared with those obtained for analogous models, like TSR or models derived from the generalized uncertainty principle. A greater variety of possibilities arises in the present case for the modifications of the uncertainty relation and for the spectrum of the harmonic oscillator, because of the freedom to choose independently the signs of the coupling constants.
The most interesting case seems to be that of negative deformation constants, since both minimal length and momentum arise. In that case, the energy of the free particle is not quantized, and the correction to the harmonic oscillator spectrum have positive sign.
Further investigations are required to understand possible phenomenological implications of these results. For example, their relevance for statistical mechanics has been studied in [7].

### v2

@@ Abstract
To give an interpretation to the physics associated to the Yang algebra, it is natural to identify $\alpha$ and $\beta$ with the cosmological constant and the inverse of the Planck mass squared, respectively. The parameters are therefore exceedingly small for experimental scales related to particle physics and gravity. Although from a mathematical point of view all values of positions and momenta are allowed (except that for positive values of the coupling constants $x^{2}<1/\alpha$ and $p^{2}<1/\beta$), the theory is supposed to be reliable only for values of $x^{2}\ll 1/\alpha$ and $p^{2}\ll 1/\beta$.
Although the Yang model has attracted growing interest recently, only its formal properties have been investigated [6], with the exception of ref. [7], where some implications on statistical mechanics and on the uncertainty principle were considered. Moreover, connections with different theories, like conformal field theory or fuzzy gravity, have been proposed by some authors [8].
@@
In this framework, we can discuss the modification of the uncertainty relation induced by the deformed symplectic structure. Deformed uncertainty relation have earned much interest in the last years, since they may emerge in a variety of situations, including string theory, quantum gravity, noncommutative geometry or gedanken experiments and are strictly related to the existence of a minimal length [9]. It must be remarked, however, that not all deformed uncertainty relations predict the existence of a minimal length [10]. As we shall see, this property is related to the sign of the correction to the Heisenberg relations.
However, also the curvature of spacetime leads to a deformation of the uncertainty relation [11]. Both deformation can be unified in the so called Extended Generalized Uncertainty Principle (EGUP) [12,13]. Of course, as we shall show, it is this last case that occurs in the Yang model.
In the context of deformed uncertainty relations, also the investigation of simple quantum mechanical models, like the free particle or the harmonic oscillator, is interesting and has been performed by several authors [14,15]. In particular, it has been shown that in some cases the free particle has discrete spectrum, while the harmonic oscillator always displays corrections with respect to the standard case. More recently, more realistic models have been studied in the context of molecular dynamics, as for example diatomic molecules [16].
@@
In the present case the investigation of the deformations of the uncertainty relation and of the dynamics can be done only if one considers a specific realization of the model on phase space, since, as noted above, it is otherwise difficult to give a physical interpretation to the scalar generator of the algebra. Moreover, to avoid computational complications, we shall limit our study to the one-dimensional case and to the leading order in $\hbar$, $\alpha$ and $\beta$. We shall consider both positive and negative values of the coupling parameters, since they lead to very different physical predictions.
@@
We shall give a heuristic derivation of our results. A rigorous mathematical treatment, including for example the discussion of the functional analysis of the position and momentum operators, would require much more details, and we plan to expose it in a forthcoming paper.
@@
${\hat{x}}_{\mu}=x_{\mu}\sqrt{\sqrt{1+\alpha\beta(x\!\cdot\!p)^{2}}-\beta p^{2}}\,\qquad{\hat{p}}_{\mu}=p_{\mu}\sqrt{\sqrt{1+\alpha\beta(x\!\cdot\!p)^{2}}-\alpha x^{2}}\ ,$ (3)3( 3 )
@@
$K=\sqrt{1-\alpha{\hat{x}}^{2}-\beta{\hat{p}}^{2}-{\alpha\beta\over 2}M_{\rho\sigma}^{2}}=\sqrt{1-\alpha{\hat{x}}^{2}-\beta{\hat{p}}^{2}-\alpha\beta\,{{\hat{x}}^{2}{\hat{p}}^{2}-({\hat{x}}\!\cdot\!{\hat{p}})^{2}\over K^{2}}}.$ (5)5( 5 )
From (3) it follows that if $\alpha\beta<0$, the scalar product $x\!\cdot\!p$ must have an upper bound $1/\sqrt{|\alpha\beta|}$, while this is not true if $\alpha\beta>0$.
@@
In the nonrelativistic model, the commutation relations are the same as (1), but now the indices run from 1 to 3, $\eta_{\mu\nu}\to\delta_{\mu\nu}$ and the algebra reduces to $o(5)$, $o(1,4)$ or $o(2,3)$, depending on the signs of $\alpha$ and $\beta$.
3. Generalized uncertainty principle A discussion of the properties of the Yang model in the realization of the previous section is however extremely difficult, even in the 3D Euclidean version, but the problem is highly simplified in the one-dimensional limit. Although in one dimension one cannot speak of noncommutativity and curvature of spacetime, this limit has nevertheless proven to maintain the most interesting features of its higher-dimensional counterparts in similar models [18], since its behaviour essentially coincides with that of the higher-dimensional radial equation, corresponding to the zero angular momentum sector. In the Snyder case this limit has also been widely investigated in the context of minimal length models [14,15]. Recently, the appearance of deformed uncertainty relation in the 1D Yang model has independently been observed in [7].
In our approximation, the one-dimensional commutation relation for ${\hat{x}}$ and ${\hat{p}}$ , with $K$ given by (5), simplifies to
$[{\hat{x}},{\hat{p}}]=i\hbar\sqrt{1-\alpha{\hat{x}}^{2}-\beta{\hat{p}}^{2}}.$ (7)7( 7 )
It follows that for $\alpha>0$, the acceptable values of ${\hat{x}}^{2}$ are bounded by $1/\alpha$, and for $\beta>0$ those of ${\hat{p}}^{2}$ by $1/\beta$, while no such bound is instead present for negative values of the coupling constants. Moreover, in general one must require $\alpha{\hat{x}}^{2}+\beta{\hat{p}}^{2}\leq 1$.
Assuming for simplicity $\langle{\hat{x}}\rangle=\langle{\hat{p}}\rangle=0$, by the Robinson-Schrödinger inequality, the commutation relation (7) implies a generalized uncertainty principle, analogous to EGUP,1 Note that in most cases studied in the literature the corrections are defined with opposite sign.
$\Delta{\hat{x}}\Delta{\hat{p}}\geq{1\over 2}\left|\langle[{\hat{x}},{\hat{p}}]\rangle\right|\approx{\hbar\over 2}\left(1-{\alpha\over 2}(\Delta{\hat{x}})^{2}-{\beta\over 2}(\Delta{\hat{p}}^{2})\right),$ (8)8( 8 )
where we have expanded at first order in the parameters $\alpha$ and $\beta$. Mathematically, the expansion is valid for ${\hat{x}}^{2}<1/\alpha$ if $\alpha>0$ and ${\hat{p}}^{2}<1/\beta$ if $\beta>0$, while it is always valid for negative values of the parameters. Moreover, considering only first-order terms is justified if ${\hat{x}}^{2}\ll 1/\alpha$ and ${\hat{p}}^{2}\ll 1/\beta$, which is always true for physically reasonable values of the coupling constants.
Taking the square of (8) gives
$(\Delta{\hat{x}})^{2}(\Delta{\hat{p}})^{2}\geq{\hbar^{2}\over 4}\Big{(}1-\alpha(\Delta{\hat{x}})^{2}-\beta(\Delta{\hat{p}})^{2}\Big{)}.$ (9)9( 9 )
@@
$(\Delta{\hat{x}})\geq\hbar\sqrt{{1-\beta(\Delta{\hat{p}})^{2}\over\hbar^{2}\alpha+4(\Delta{\hat{p}})^{2}}},\qquad(\Delta{\hat{p}})\geq\hbar\sqrt{{1-\alpha(\Delta{\hat{x}})^{2}\over\hbar^{2}\beta+4(\Delta{\hat{x}})^{2}}}.$ (10)10( 10 )
The behaviour of $\Delta x$ in function of $\Delta p$ in the various cases is reported in fig. 1.
For $\alpha,\beta>0$, the uncertainties satisfy the bounds $0\leq\Delta{\hat{p}}\leq 1/\sqrt{\beta}$, and $0\leq\Delta{\hat{x}}\leq 1/\sqrt{\alpha}$, as is also obvious because of the bounds imposed on ${\hat{x}}$ and ${\hat{p}}$.
For $\alpha,\beta<0$, $\Delta{\hat{p}}\geq\hbar\sqrt{|\alpha|}/2$, and $\Delta{\hat{x}}\geq\hbar\sqrt{|\beta|}/2$, hence both position and momentum have a minimal uncertainty independent of that of the conjugated variable.
A more complicated behaviour occurs if $\alpha$ and $\beta$ have different signs. If $\alpha<0$ and $\beta>0$, $\Delta{\hat{p}}$ must satisfy the bounds $\hbar\sqrt{|\alpha|}/2\geq\Delta{\hat{p}}\geq 1/\sqrt{\beta}$, while $\Delta{\hat{x}}$ can take any positive value.
Finally, if $\alpha>0$ and $\beta<0$, $\Delta p$ can take any positive value, while $\hbar\sqrt{|\beta|}/2\geq\Delta x\geq 1/\sqrt{\alpha}$. It is easy to check that in all cases the minimal values so obtained satisfy the bound $1-\alpha{\hat{x}}^{2}-\beta{\hat{p}}^{2}\geq 0$. It is also clear that if either $\alpha$ or $\beta$ is positive, it is possible to have states with vanishing $\Delta{\hat{x}}\Delta{\hat{p}}$, while this cannot occur if both $\alpha$ and $\beta$ are negative. It seems therefore that only the case of negative $\alpha$ and $\beta$ can satisfy the standard uncertainty bound. However, also for positive values of the coupling constant states with vanishing uncertainty are achieved only for values of $\Delta{\hat{x}}$ and $\Delta{\hat{p}}$ close to the limit values of the phase space variables, which are assumed to be out of the range of applicability of the theory.
Of course, our derivation has been heuristic. A rigorous discussion would require a definition of squeezed states on a suitable Hilbert space, in order to define maximally localized states [14,21,12]. This will be discussed in detail in a forthcoming paper.
@@
$\Delta{\hat{x}}\Delta{\hat{p}}\geq{\hbar\over 2}\,{1+\alpha\Delta{\hat{x}}^{2}+\beta\Delta{\hat{p}}^{2}\over 1+\hbar\sqrt{\alpha\beta}}.$ (11)11( 11 )
@@
$(\Delta{\hat{x}}^{2})_{min}={\hbar^{2}\beta\over 1+2\hbar\sqrt{\alpha\beta}},\qquad(\Delta{\hat{p}})^{2}_{min}={\hbar^{2}\alpha\over 1+2\hbar\sqrt{\alpha\beta}}.$ (12)12( 12 )
@@ 4. One-dimensional dynamics
Clearly, the quantum dynamics of the Yang model differs from the standard one, because of the deformation of the commutation relations. In one dimension, the dynamics of the free particle and of the harmonic oscillator has been studied for the Snyder model (also in the context of the generalized uncertainty principle) [14,15,5] and TSR [18], displaying a discrete spectrum for the free particle with positive coupling constants, and a deformation of the energy spectrum of the harmonic oscillator.
Let us then consider the dynamics of these models in the Yang case. For simplicity, we shall limit our study to the case of positive $\alpha$ and $\beta$. We recall that in this case $|x|<1/\alpha$, $|p|<1/\beta$, but no minimal uncertainties occur. We start by considering the solution of the one-dimensional Schrödinger equation for a free particle, using the realization (3).
Since the equation obtained in this way is not exactly solvable, we make an expansion in the parameters $\alpha$ and $\beta$. In one dimension, eqs. (3) become at leading order
${\hat{x}}_{\mu}=x_{\mu}-{\beta\over 4}(x_{\mu}p^{2}+p^{2}x_{\mu}),\qquad{\hat{p}}_{\mu}=p_{\mu}-{\alpha\over 4}(p_{\mu}x^{2}+x^{2}p_{\mu}),$ (13)13( 13 )
where we have chosen a Hermitian realization.
One can now work either in position or momentum representation. In the first case,
$x_{\mu}\to x_{\mu},\quad p_{\mu}\to-i{\partial\over\partial x_{\mu}}.$ (14)14( 14 )
Although the choice of a Hilbert space for the exact theory is difficult, in the context of our approximation we may take it as the space of square-integrable functions of $x$ in the interval $|x|\leq 1/\sqrt{\alpha}$, with standard scalar product. At leading order this is consistent with the requirement that $|{\hat{x}}|\leq 1/\sqrt{\alpha}$, which is natural in our representation.
The one-dimensional Schrödinger equation for a free particle thus becomes
$({\hat{p}}^{2}-2mE)\,\psi=-(1-\alpha x^{2})\,{d^{2}\psi\over dx^{2}}+2\alpha x{d\psi\over dx}+\left({\alpha\over 2}-2mE\right)\psi=0.$ (15)15( 15 )
Requiring that $\psi$ and its derivative are regular at $x=\pm 1/\sqrt{\alpha}$, the solution is given by
$\psi\propto P_{n}(\sqrt{\alpha}\,x),$ (16)16( 16 )
@@
$E_{n}={\alpha\over 2m}\left(n^{2}+n+{1\over 2}\right).$ (17)17( 17 )
@@
$p_{\mu}\to p_{\mu},\quad x_{\mu}\to i{\partial\over\partial p_{\mu}},$ (18)18( 18 )
@@
$({\hat{p}}^{2}-2mE)\,\psi=\alpha p^{2}{d^{2}\psi\over dp^{2}}+2\alpha p\,{d\psi\over dp}+\left(p^{2}-2mE+{\alpha\over 2}\right)\psi=0.$ (19)19( 19 )
The regular solution is given by
$\psi\propto j_{n}\left({p\over\sqrt{\alpha}}\right),$ (20)20( 20 )
with $j_{n}$ a spherical Bessel function with index $n$ such that $n(n+1)={2mE\over\alpha}-{1\over 2}$, recovering the previous result.
@@
$H={{\hat{p}}^{2}\over 2m}+{m\omega^{2}{\hat{x}}^{2}\over 2}.$ (21)21( 21 )
@@
$\Big{(}1-A^{2}x^{2}\Big{)}\,{d^{2}\psi\over dx^{2}}-2A^{2}x{d\psi\over dx}+\left(2mE-{A^{2}\over 2}-m^{2}\omega^{2}x^{2}\right)\psi=0,$ (22)22( 22 )
@@
$A^{2}=\alpha+\beta m^{2}\omega^{2}.$ (23)23( 23 )
@@
$(1-z^{2})\,{d^{2}\psi\over dz^{2}}-2z{d\psi\over dz}+\left({2mE\over A^{2}}-{1\over 2}-{m^{2}\omega^{2}\over A^{4}}z^{2}\right)\psi=0,$ (24)24( 24 )
@@
$\psi_{n}\propto S_{0n}(\gamma,Ax),$ (25)25( 25 )
with $\gamma={m\omega\over A^{2}}$. Notice that eq. (24) has not a smooth limit for $A\to 0$. From (24) it follows that
$E_{n}={A^{2}\over 2m}\left(\lambda_{n}+{1\over 2}\right).$ (26)26( 26 )
@@
$\lambda_{n}=\sum_{k=0}^{\infty}l_{2k}\gamma^{2k},$ (27)27( 27 )
@@
$l_{0}=n(n+1),\qquad l_{2}={1\over 2}\left(1-{1\over(2n-1)(2n+3)}\right),\qquad{\rm etc.}$ (28)28( 28 )
@@
$\lambda_{n}=2\left(n+{1\over 2}\right)\gamma-{1\over 2}\left(n^{2}+n+{3\over 2}\right)+o\left({1\over\gamma}\right),$ (29)29( 29 )
@@
$E_{n}=\left(n+{1\over 2}\right)\omega-{A^{2}\over 4m}\left(n^{2}+n+{1\over 2}\right)+o\left({A^{4}\over m^{2}\omega}\right).$ (30)30( 30 )
Besides the standard term, one has corrections proportional to powers of $A^{2}$. Also this result is analogous to TSR [18], where $E_{n}=\left(n+{1\over 2}\right)\omega+{A^{2}\over 2m}\left(n^{2}+n+{1\over 2}\right)+o\left({A^{4}\over m^{2}\omega}\right)$, but the coefficients of the correction terms are different, and in particular have opposite sign.
@@
5. Final remarks We have considered the quantum mechanics of the one-dimensional Yang model in its simplest realization on phase space. The investigation of this simple case is relevant because, as shown in similar situations [15,18], the results are qualitatively similar to those valid in three dimensions. The results can be compared with those obtained for analogous models, like TSR or models derived from the generalized uncertainty principle. A greater variety of possibilities arises in the present case for the modifications of the uncertainty relation and for the spectrum of the harmonic oscillator, because of the freedom to choose independently the signs of the coupling constants. As already shown in [10] and [5,18], in fact, changing the signs of the couplings in the deformed uncertainty relation greatly modifies the physics, leading for example to the disappearance of the minimal length.
From the results obtained, the most interesting case in our model seems to be that of negative deformation constants, since both minimal length and momentum arise. In that case, the energy of the free particle is not quantized, and the correction to the harmonic oscillator spectrum have positive sign.
In this paper we have not been mathematically rigorous. A more thorough investigation would require the definition of a suitable Hilbert space and the study of the functional analysis of the position and momentum operators, in particular when minimal uncertainties are present, like in the similar case of two-parameter deformation of the uncertainty principle studied in [12]. This analysis is more involved and we plan to afford it in a forthcoming paper.
Further investigations are also required to understand possible phenomenological implications of our results. For example, their relevance for statistical mechanics has been studied in [7]. Also applications to molecular dynamics in analogy with ref. [16] can be investigated.